OpenStudy (anonymous):

Part 1: Solve the inequality −5x + 8 x − 4 and show all of your work. You can use <= to represent the less than or equal to symbol. Part 2: Use complete sentences to describe the graph of the solution from Part 1. Be sure to include the type of circle and which direction the shading will be.

OpenStudy (anonymous):

\[-5x + 8 \le x-4\]

OpenStudy (anonymous):

thats part 1

OpenStudy (ganpat):

try to isolate x... can u do that ??

OpenStudy (anonymous):

i'm not sure

OpenStudy (ganpat):

bring x on one side and all values on other..

OpenStudy (ganpat):

−5x+8 ≤ x−4 5x + x >= 8 + 4 6x >= 12 x >= 2.. something like this ??

OpenStudy (ganpat):